OpenStudy (anonymous):

The volume of a pyramid that fits exactly inside a cube is 10 cubic feet. What is the volume of the cube? (5 points) Select one: a. 30 cubic feet b. 100 cubic feet c. 300 cubic feet d. 1000 cubic feet

OpenStudy (anonymous):

@amistre64 @AravindG @agent0smith @arabpride @arabpride @atlas @aprilann143 @Barrelracer011 @brei @baby456 @BrianMatthew @cookiibabii93 @ctfxcfan1 @Cjsmall14 @Destinymasha @DollyAcquah @Darkesage @Dodo1 @ess=bae @f

OpenStudy (anonymous):

I just need help like understanding this

OpenStudy (anonymous):

and I'm not sure how to do this

OpenStudy (anonymous):

wat u want

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

what do you think lol?

OpenStudy (anonymous):

IDK u mentioned me so im confuzzesd as f so plllzzzzzzzzz..... inlighten me

OpenStudy (anonymous):

bro leave lol

OpenStudy (anonymous):

i get it if you don't understand lol sorry I asked you I wasn't sure buddy

OpenStudy (anonymous):

what do you understand and ill help from there

OpenStudy (anonymous):

okay i understand I need to find the volume

OpenStudy (anonymous):

okay 1 min

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ou can put this solution on YOUR website! Volume of Pyramid = (1/3)*(Volume of Prism) 3*(Volume of Pyramid) = Volume of Prism Volume of Prism = 3*(Volume of Pyramid)

OpenStudy (anonymous):

Volume of Prism = 3*(10) Volume of Prism = 30 The volume of the cube (which is a special type of rectangular prism) is 30 cubic feet

OpenStudy (anonymous):

So what is your answer ?

OpenStudy (anonymous):

I think its 30?

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

@ryan>>>

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

np

OpenStudy (calculusfunctions):

The volume of any pyramid is always one-third the volume of the prism, in which the pyramid exactly fits.